OpenStudy (anonymous):

differentiate x^(-1) twice....please

OpenStudy (anonymous):

\[d ^{2}/dx ^{2}(1/x) \implies d/dx(-1/x ^{2})\implies 2/x ^{3}\]

OpenStudy (anonymous):

\[d/dx(x ^{n})=nx ^{n-1}\]

OpenStudy (anonymous):

n is any integer

OpenStudy (anonymous):

ok, and just regular x would first become x^(-1), then -x^(-2)?

OpenStudy (lgbasallote):

\[\frac{d}{dx} x^{-1} \implies -1(x)^{-1 - 1} \implies -x^{-2}\] \[\frac{d}{dx} x^{-2} \implies -2 (x)^{-2 - 1} \implies -2x^{-3}\]

OpenStudy (lgbasallote):

sorry that should be positive

OpenStudy (anonymous):

thank you